OpenStudy (anonymous):

find the slope, if it exists 7x+6y=42

OpenStudy (anonymous):

put this in the form y=mx+b. then solve for y. the number in front of x is your slope. so 6y=-7x=42. then y=-7/6x=7. -7/6 is your slope or "m".

OpenStudy (anonymous):

"6y=-7x=42" shoud be "6y=-7x+42" whoops